# An Evidence-Based Post-Hoc Adjustment Framework for Anomaly Detection Under Data Contamination

### v1

@@ Abstract
Unsupervised anomaly detection (AD) methods typically assume clean training data, yet real-world datasets often contain undetected or mislabeled anomalies, leading to significant performance degradation. Existing solutions require access to the training pipelines, data or prior knowledge of the proportions of anomalies in the data, limiting their real-world applicability. To address this challenge, we propose EPHAD, a simple yet effective test-time adaptation framework that updates the outputs of AD models trained on contaminated datasets using evidence gathered at test time. Our approach integrates the prior knowledge captured by the AD model trained on contaminated datasets with evidence derived from multimodal foundation models like Contrastive Language-Image Pre-training (CLIP), classical AD methods like the Latent Outlier Factor or domain-specific knowledge. We illustrate the intuition behind EPHAD using a synthetic toy example and validate its effectiveness through comprehensive experiments across eight visual AD datasets, twenty-six tabular AD datasets, and a real-world industrial AD dataset. Additionally, we conduct an ablation study to analyse hyperparameter influence and robustness to varying contamination levels, demonstrating the versatility and robustness of EPHAD across diverse AD models and evidence pairs. To ensure reproducibility, our code is publicly available [^1] https://github.com/sukanyapatra1997/EPHAD.
@@ 1 Introduction
Anomaly detection (AD) is the basis of many critical applications, including cybersecurity (Xiao et al.,, 2024; Li et al., 2023a, ), healthcare (Bijlani et al.,, 2024; Huang et al.,, 2024), and industrial maintenance (Schwarz et al.,, 2025; Patra et al.,, 2024). By enabling the identification of abnormalities, potential threats, or critical system failures, AD contributes to the robustness and safety of real-world systems. Despite its significance, AD remains a challenging task due to the inherent difficulty in characterising anomalous behaviours and the lack of prior knowledge about anomalous samples (Ruff et al.,, 2021). Consequently, AD is commonly approached as an unsupervised representation learning problem without access to labelled anomalies (Batzner et al.,, 2024; You et al.,, 2022).
@@
A standard approach in unsupervised AD involves training a model to learn a “compact” representation of the normal samples from a training dataset under the assumption that the training data is “clean", i.e. contains only normal samples (Ruff et al.,, 2021). Then, anomalies are identified as deviations from this learned normality. One-class (OC) classification methods (Ruff et al.,, 2018; Tax and Duin,, 2004) learn a decision boundary that encompasses all the normal samples. In contrast, density-based methods (Gudovskiy et al.,, 2022; Yu et al.,, 2021) learn the probability distribution of normal samples. Furthermore, memory bank-based approaches (Roth et al.,, 2022) store the features corresponding to normal samples in a memory bank. However, real-world datasets are often contaminated with undetected anomalies (Das et al.,, 2025; Hien et al.,, 2024; Qiu et al.,, 2022). For example, a dataset collected for industrial maintenance may already include unnoticed defects. This leads to biased AD models that struggle to reliably distinguish between normal and anomalous instances.
@@
We consider the more realistic setting where the training data may be contaminated with anomalies. Existing approaches to handle contamination in the unsupervised setting primarily follow two strategies. The first employs an auxiliary OC classifier to filter out suspected anomalies (Yoon et al.,, 2022; Jiang et al.,, 2022), while the second modifies the training pipeline to enhance robustness against contamination (Qiu et al.,, 2022; Eduardo et al.,, 2020). Although effective, these methods rely on prior knowledge of the proportion of anomalies in the training data, i.e. the contamination ratio, which is typically unknown. Also, such methods are often computationally expensive. In the semi-supervised setting, methods leverage additional labelled datasets containing normal and anomalous samples (Hien et al.,, 2024; Ruff et al.,, 2020). However, their effectiveness diminishes when the anomalous instances encountered during training do not replicate real-world anomalies (Perini et al.,, 2025).
@@
In this work, we aim to mitigate the possible adverse effects of data contamination on the performance of unsupervised AD models (Bouman et al.,, 2024). Specifically, we address the challenging setting in which training pipelines, data, or prior knowledge of the proportions of anomalies cannot be accessed. This scenario reflects the growing trend of deploying proprietary AD models in real-world applications, where access to internal model components is often restricted. Even when fine-tuning is permitted, it is not only computationally intensive but also unreliable due to the absence of guaranteed clean training data, as anomalies are inherently unknown a priori. This setup aligns with preparation-agnostic test-time adaptation (TTA) methods (Karmanov et al.,, 2024; Zhang et al.,, 2023; Xiao and Snoek,, 2024) , which remain largely unexplored in the context of AD. To address this gap, we introduce the Evidence-based Post-Hoc Adjustment Framework for Anomaly Detection (EPHAD), a simple yet effective method that adjusts the outputs of a pretrained AD model post-hoc, using evidence collected at test time.
@@
Notably, we establish conceptual links between EPHAD and recent advances in test-time alignment for generative models (Mudgal et al.,, 2024; Li et al.,, 2024; Korbak et al.,, 2022), underscoring its broader significance. EPHAD is flexible and can incorporate various forms of evidence, including foundation models like Contrastive Language–Image Pre-training (CLIP) (Zhou et al.,, 2024; Jeong et al.,, 2023), classical AD methods such as Latent Outlier Factor (LOF) (Breunig et al.,, 2000), and domain-specific knowledge. Our core contributions are summarised below:
@@ 2 Related work
Unsupervised AD. Over the years, numerous approaches have been developed for unsupervised AD, which can be broadly categorised into four main families: one-class classifiers (OCCs), feature embedding-based, density-based, and reconstruction-based methods. OCCs aim to learn a decision boundary that encapsulates all normal samples. Classical OCC approaches employ shallow models such as support vector-based methods that learn a maximum-margin hyperplane (Schölkopf et al.,, 2001) or a hypersphere (Tax and Duin,, 1999). To mitigate the limitations of manual feature engineering and extend to high-dimensional data, deep learning-based variants like DeepSVDD (Ruff et al.,, 2018) have been introduced. Feature embedding-based methods, on the other hand, leverage pre-trained deep models to extract representations of input data. These representations are then either stored in a memory bank (Roth et al.,, 2022; Lee et al.,, 2022) or used to train a student-teacher network (Zhang et al.,, 2024; Batzner et al.,, 2024; Patra and Ben Taieb,, 2024). Density-based methods detect anomalies by estimating the probability distribution of normal samples, assuming that anomalies reside in low-density regions. While early methods include KDE (Kim and Scott,, 2012), more recent deep-learning-based variants include DAGMM (Zong et al.,, 2018), CFLOW (Gudovskiy et al.,, 2022), and FastFlow (Yu et al.,, 2021). Lastly, reconstruction-based approaches learn to map normal samples into a lower-dimensional bottleneck and reconstruct them. The inability to accurately reconstruct samples during inference serves as a detection criterion. For a more comprehensive survey, we refer readers to Liu et al., (2024) and Ruff et al., (2021).
@@
Data contamination. Handling dataset contamination in AD typically assumes a low proportion of anomalies, allowing methods to prioritise normal instances (inlier priority) (Wang et al.,, 2019). However, in practice, this assumption is difficult to ensure since anomalies are often unknown. To mitigate contamination, Yoon et al., (2022) proposed a data refinement approach using an ensemble of one-class classifiers (OCCs) to filter suspected anomalies and create a cleaner dataset. While effective, this method incurs high computational costs and discards anomalies rather than leveraging them for improved generalisation via Outlier Exposure (Hendrycks et al.,, 2019). To address this, Qiu et al., (2022) introduced Latent Outlier Exposure (LOE), which iteratively assigns anomaly scores and infers labels using block coordinate descent while incorporating the contamination ratio to prevent degenerate solutions. However, estimating the contamination ratio remains a challenge. Perini et al., (2022) tackled this by leveraging an auxiliary dataset with a known contamination ratio, assuming domain similarity. Alternatively, Perini et al., (2023) fits a Dirichlet Process Gaussian Mixture Model to anomaly scores, though this approach lacks a closed-form solution. Despite these advancements, existing methods introduce computational overhead and are often impractical for modern pre-trained proprietary models, limiting their real-world applicability.
@@ 3 Background
Density-based anomaly detection. An anomaly can be defined as “an observation that deviates significantly from some concept of normality” (Ruff et al.,, 2021). This definition comprises two key aspects: the concept of normality and the significant deviation from it, which can be formalised using a probabilistic framework. The concept of normality is defined as the probability distribution of normal samples $\textup{P}^{+}_{X}$. To formalise this further, we adopt the concentration assumption (Steinwart et al.,, 2005), which posits that although the data space $\mathcal{X}$} is unbounded, the high-density regions of $\textup{P}^{+}_{X}$ are bounded and concentrated. In contrast, $\textup{P}^{-}_{X}$ is assumed to be non-concentrated (Schölkopf and Smola,, 2002), and is often approximated by a uniform distribution over $\mathcal{X}$} (Tax,, 2001). Given the PDF $f^{+}_{X}$ associated with $\textup{P}^{+}_{X}$, which we refer to as inlier density, a data point $x\in{\mathcal{X}}$ is identified as an anomaly if it deviates substantially from this concept of normality, i.e., if it resides in a low-probability region under $\textup{P}^{+}_{X}$. However, since $f^{+}_{X}$ is typically unknown in practice, density-based methods approximate it using a density estimator.
Score-based anomaly detection. Density estimation poses significant challenges, particularly in high-dimensional spaces or when data is sparse, and often incurs substantial computational cost. Fortunately, in the context of anomaly detection, the goal is typically not to recover the exact data likelihood but rather to establish a ranking of data points based on their degree of normality. This motivates an alternative strategy: learning an anomaly score function $s_{\text{out}}(x):{\mathcal{X}}\to\mathbb{R}$ , which directly assigns an anomaly score to a data point $x\in{\mathcal{X}}$, thereby quantifying its degree of anomalousness (Ruff et al.,, 2021). To complement this, the inlier score function is defined as $s_{\text{in}}(x)=-s_{\text{out}}(x)$ , capturing the degree of normality, where higher values indicate that $x$ is normal. For AD, first, we train a model to learn the anomaly score function $s^{+}_{\text{out}}(x)$ using $\mathcal{D}^{+}_{\mathrm{train}}$. Then, we define the anomaly detector as
@@
where $\lambda_{s}\geq 0$ is a pre-determined threshold (Perini et al.,, 2023, 2022). The density-based AD method can also be interpreted as a specific case of the score-based AD methods where the anomaly score $s^{+}_{\text{out}}(x)=-\phi(f^{+}_{X}(x))$ and $\phi(\cdot)$ is an order-preserving transformation chosen to be the logarithm.
Data contamination. For training the AD method, a common assumption is that the training dataset $\mathcal{D}^{+}_{\mathrm{train}}$ consists solely of i.i.d. samples from the normal data distribution $\textup{P}^{+}_{X}$, without anomalies. However, this assumption is rarely satisfied in practice, since anomalies are typically unknown a priori. As a result, the training dataset is often contaminated with undetected anomalies. A more realistic assumption is that the dataset $\mathcal{D}^{\pm}_{\mathrm{train}}:=\{x_{i}\}_{i=1}^{m}$ contains both normal and anomalous samples drawn from a mixture distribution $\textup{P}^{\pm}_{X}$ with PDF $f^{\pm}_{X}$ (Huber and Ronchetti,, 2011; Huber,, 1992). Letting $\epsilon=\mathbb{P}(Y=-1)$ denote the contamination factor, $\textup{P}^{\pm}_{X}$ can be written as
@@
As $\epsilon$ increases, the model trained on $\mathcal{D}^{\pm}_{\mathrm{train}}$ becomes biased towards the anomalous regions, reducing its ability to separate normal from anomalous samples (Qiu et al.,, 2022; Yoon et al.,, 2022). The existing literature examining the impact of contamination on unsupervised AD methods (Jiang et al.,, 2022; Qiu et al.,, 2022; Hien et al.,, 2024; Perini et al.,, 2023, 2022) typically considers contamination levels ranging from $0\%$ to $20\%$. Additionally, an analysis of 57 datasets spanning Natural Language Processing and Computer Vision in ADBench (Han et al.,, 2022) [Appendix B.2, Figure B1] revealed that nearly $70\%$ of the datasets exhibit anomaly ratios below 10%, with a median of $5\%$.
@@ 4 EPHAD: An evidence-based post-hoc adjustment framework
We consider the realistic scenario in which an AD model has already been trained on a possibly contaminated dataset $\mathcal{D}^{\pm}_{\mathrm{train}}$. Instead of retraining the model, our goal is to adapt its test-time predictions to mitigate the impact of contamination. To this end, we introduce a novel Evidence-based Post-Hoc Adjustment Framework for Anomaly Detection (EPHAD), that corrects model predictions using an evidence function at test-time. The evidence function $T(x):\mathcal{X}\to\mathbb{R}$ assigns higher values to samples deemed more likely to be normal and can incorporate domain-specific knowledge. Thus, EPHAD aligns with preparation-agnostic TTA methods (Xiao and Snoek,, 2024).
@@
This objective balances two competing goals: aligning the adjusted PDF with the evidence (first term) and maintaining fidelity to the original PDF (second term). The temperature parameter $\beta$ controls this trade-off, recovering the evidence-driven solution as $\beta\to 0$ and reverting to the original model as $\beta\to\infty$. For the proof of (5), see Korbak et al., (2022). This interpretation also highlights a close connection to well-established TTA approaches used in generative models (Korbak et al.,, 2022; Mudgal et al.,, 2024; Li et al.,, 2024), where the model is viewed as an RL policy fine-tuned with a reward function encoding evidence or alignment criteria. In this view, EPHAD performs a KL-regularised shift of the contaminated density $f^{\pm}_{X}$ toward regions favored by the evidence function $T(x)$ while preserving consistency with $f^{\pm}_{X}$ through the KL term.

### v2

@@ Abstract
Unsupervised anomaly detection (AD) methods typically assume clean training data, yet real-world datasets often contain undetected or mislabeled anomalies, leading to significant performance degradation. Existing solutions require access to the training pipelines, data or prior knowledge of the proportions of anomalies in the data, limiting their real-world applicability. To address this challenge, we propose EPHAD, a simple yet effective test-time adaptation framework that updates the outputs of AD models trained on contaminated datasets using evidence gathered at test time. Our approach integrates the prior knowledge captured by the AD model trained on contaminated datasets with evidence derived from multimodal foundation models like Contrastive Language-Image Pre-training (CLIP), classical AD methods like the Local Outlier Factor or domain-specific knowledge. We illustrate the intuition behind EPHAD using a synthetic toy example and validate its effectiveness through comprehensive experiments across eight visual AD datasets, twenty-six tabular AD datasets, and a real-world industrial AD dataset. Additionally, we conduct an ablation study to analyse hyperparameter influence and robustness to varying contamination levels, demonstrating the versatility and robustness of EPHAD across diverse AD models and evidence pairs. To ensure reproducibility, our code is publicly available [^1] https://github.com/sukanyapatra1997/EPHAD.
@@ 1 Introduction
Anomaly detection (AD) is the basis of many critical applications, including cybersecurity (Xiao et al., 2024; Li et al., 2023a), healthcare (Bijlani et al., 2024; Huang et al., 2024), and industrial maintenance (Schwarz et al., 2025; Patra et al., 2024). By enabling the identification of abnormalities, potential threats, or critical system failures, AD contributes to the robustness and safety of real-world systems. Despite its significance, AD remains a challenging task due to the inherent difficulty in characterising anomalous behaviours and the lack of prior knowledge about anomalous samples (Ruff et al., 2021). Consequently, AD is commonly approached as an unsupervised representation learning problem without access to labelled anomalies (Batzner et al., 2024; You et al., 2022).
@@
A standard approach in unsupervised AD involves training a model to learn a “compact” representation of the normal samples from a training dataset under the assumption that the training data is “clean", i.e. contains only normal samples (Ruff et al., 2021). Then, anomalies are identified as deviations from this learned normality. One-class (OC) classification methods (Ruff et al., 2018; Tax and Duin, 2004) learn a decision boundary that encompasses all the normal samples. In contrast, density-based methods (Gudovskiy et al., 2022; Yu et al., 2021) learn the probability distribution of normal samples. Furthermore, memory bank-based approaches (Roth et al., 2022) store the features corresponding to normal samples in a memory bank. However, real-world datasets are often contaminated with undetected anomalies (Das et al., 2025; Hien et al., 2024; Qiu et al., 2022). For example, a dataset collected for industrial maintenance may already include unnoticed defects. This leads to biased AD models that struggle to reliably distinguish between normal and anomalous instances.
@@
We consider the more realistic setting where the training data may be contaminated with anomalies. Existing approaches to handle contamination in the unsupervised setting primarily follow two strategies. The first employs an auxiliary OC classifier to filter out suspected anomalies (Yoon et al., 2022; Jiang et al., 2022), while the second modifies the training pipeline to enhance robustness against contamination (Qiu et al., 2022; Eduardo et al., 2020). Although effective, these methods rely on prior knowledge of the proportion of anomalies in the training data, i.e. the contamination ratio, which is typically unknown. Also, such methods are often computationally expensive. In the semi-supervised setting, methods leverage additional labelled datasets containing normal and anomalous samples (Hien et al., 2024; Ruff et al., 2020). However, their effectiveness diminishes when the anomalous instances encountered during training do not replicate real-world anomalies (Perini et al., 2025).
@@
In this work, we aim to mitigate the possible adverse effects of data contamination on the performance of unsupervised AD models (Bouman et al., 2024). Specifically, we address the challenging setting in which training pipelines, data, or prior knowledge of the proportions of anomalies cannot be accessed. This scenario reflects the growing trend of deploying proprietary AD models in real-world applications, where access to internal model components is often restricted. Even when fine-tuning is permitted, it is not only computationally intensive but also unreliable due to the absence of guaranteed clean training data, as anomalies are inherently unknown a priori. This setup aligns with preparation-agnostic test-time adaptation (TTA) methods (Karmanov et al., 2024; Zhang et al., 2023; Xiao and Snoek, 2024) , which remain largely unexplored in the context of AD. To address this gap, we introduce the Evidence-based Post-Hoc Adjustment Framework for Anomaly Detection (EPHAD), a simple yet effective method that adjusts the outputs of a pretrained AD model post-hoc, using evidence collected at test time.
@@
Notably, we establish conceptual links between EPHAD and recent advances in test-time alignment for generative models (Mudgal et al., 2024; Li et al., 2024; Korbak et al., 2022), underscoring its broader significance. EPHAD is flexible and can incorporate various forms of evidence, including foundation models like Contrastive Language–Image Pre-training (CLIP) (Zhou et al., 2024; Jeong et al., 2023), classical AD methods such as Local Outlier Factor (LOF) (Breunig et al., 2000), and domain-specific knowledge. Our core contributions are summarised below:
@@ 2 Related work
Unsupervised AD. Over the years, numerous approaches have been developed for unsupervised AD, which can be broadly categorised into four main families: one-class classifiers (OCCs), feature embedding-based, density-based, and reconstruction-based methods. OCCs aim to learn a decision boundary that encapsulates all normal samples. Classical OCC approaches employ shallow models such as support vector-based methods that learn a maximum-margin hyperplane (Schölkopf et al., 2001) or a hypersphere (Tax and Duin, 1999). To mitigate the limitations of manual feature engineering and extend to high-dimensional data, deep learning-based variants like DeepSVDD (Ruff et al., 2018) have been introduced. Feature embedding-based methods, on the other hand, leverage pre-trained deep models to extract representations of input data. These representations are then either stored in a memory bank (Roth et al., 2022; Lee et al., 2022) or used to train a student-teacher network (Zhang et al., 2024; Batzner et al., 2024; Patra and Ben Taieb, 2024). Density-based methods detect anomalies by estimating the probability distribution of normal samples, assuming that anomalies reside in low-density regions. While early methods include KDE (Kim and Scott, 2012), more recent deep-learning-based variants include DAGMM (Zong et al., 2018), CFLOW (Gudovskiy et al., 2022), and FastFlow (Yu et al., 2021). Lastly, reconstruction-based approaches learn to map normal samples into a lower-dimensional bottleneck and reconstruct them. The inability to accurately reconstruct samples during inference serves as a detection criterion. For a more comprehensive survey, we refer readers to Liu et al. (2024) and Ruff et al. (2021).
@@
Data contamination. Handling dataset contamination in AD typically assumes a low proportion of anomalies, allowing methods to prioritise normal instances (inlier priority) (Wang et al., 2019). However, in practice, this assumption is difficult to ensure since anomalies are often unknown. To mitigate contamination, Yoon et al. (2022) proposed a data refinement approach using an ensemble of one-class classifiers (OCCs) to filter suspected anomalies and create a cleaner dataset. While effective, this method incurs high computational costs and discards anomalies rather than leveraging them for improved generalisation via Outlier Exposure (Hendrycks et al., 2019). To address this, Qiu et al. (2022) introduced Latent Outlier Exposure (LOE), which iteratively assigns anomaly scores and infers labels using block coordinate descent while incorporating the contamination ratio to prevent degenerate solutions. However, estimating the contamination ratio remains a challenge. Perini et al. (2022) tackled this by leveraging an auxiliary dataset with a known contamination ratio, assuming domain similarity. Alternatively, Perini et al. (2023) fits a Dirichlet Process Gaussian Mixture Model to anomaly scores, though this approach lacks a closed-form solution. Despite these advancements, existing methods introduce computational overhead and are often impractical for modern pre-trained proprietary models, limiting their real-world applicability.
@@ 3 Background
Density-based anomaly detection. An anomaly can be defined as “an observation that deviates significantly from some concept of normality” (Ruff et al., 2021). This definition comprises two key aspects: the concept of normality and the significant deviation from it, which can be formalised using a probabilistic framework. The concept of normality is defined as the probability distribution of normal samples $\textup{P}^{+}_{X}$. To formalise this further, we adopt the concentration assumption (Steinwart et al., 2005), which posits that although the data space $\mathcal{X}$} is unbounded, the high-density regions of $\textup{P}^{+}_{X}$ are bounded and concentrated. In contrast, $\textup{P}^{-}_{X}$ is assumed to be non-concentrated (Schölkopf and Smola, 2002), and is often approximated by a uniform distribution over $\mathcal{X}$} (Tax, 2001). Given the PDF $f^{+}_{X}$ associated with $\textup{P}^{+}_{X}$, which we refer to as inlier density, a data point $x\in{\mathcal{X}}$ is identified as an anomaly if it deviates substantially from this concept of normality, i.e., if it resides in a low-probability region under $\textup{P}^{+}_{X}$. However, since $f^{+}_{X}$ is typically unknown in practice, density-based methods approximate it using a density estimator.
Score-based anomaly detection. Density estimation poses significant challenges, particularly in high-dimensional spaces or when data is sparse, and often incurs substantial computational cost. Fortunately, in the context of anomaly detection, the goal is typically not to recover the exact data likelihood but rather to establish a ranking of data points based on their degree of normality. This motivates an alternative strategy: learning an anomaly score function $s_{\text{out}}(x):{\mathcal{X}}\to\mathbb{R}$ , which directly assigns an anomaly score to a data point $x\in{\mathcal{X}}$, thereby quantifying its degree of anomalousness (Ruff et al., 2021). To complement this, the inlier score function is defined as $s_{\text{in}}(x)=-s_{\text{out}}(x)$ , capturing the degree of normality, where higher values indicate that $x$ is normal. For AD, first, we train a model to learn the anomaly score function $s^{+}_{\text{out}}(x)$ using $\mathcal{D}^{+}_{\mathrm{train}}$. Then, we define the anomaly detector as
@@
where $\lambda_{s}\geq 0$ is a pre-determined threshold (Perini et al., 2023, 2022). The density-based AD method can also be interpreted as a specific case of the score-based AD methods where the anomaly score $s^{+}_{\text{out}}(x)=-\phi(f^{+}_{X}(x))$ and $\phi(\cdot)$ is an order-preserving transformation chosen to be the logarithm.
Data contamination. For training the AD method, a common assumption is that the training dataset $\mathcal{D}^{+}_{\mathrm{train}}$ consists solely of i.i.d. samples from the normal data distribution $\textup{P}^{+}_{X}$, without anomalies. However, this assumption is rarely satisfied in practice, since anomalies are typically unknown a priori. As a result, the training dataset is often contaminated with undetected anomalies. A more realistic assumption is that the dataset $\mathcal{D}^{\pm}_{\mathrm{train}}:=\{x_{i}\}_{i=1}^{m}$ contains both normal and anomalous samples drawn from a mixture distribution $\textup{P}^{\pm}_{X}$ with PDF $f^{\pm}_{X}$ (Huber and Ronchetti, 2011; Huber, 1992). Letting $\epsilon=\mathbb{P}(Y=-1)$ denote the contamination factor, $\textup{P}^{\pm}_{X}$ can be written as
@@
As $\epsilon$ increases, the model trained on $\mathcal{D}^{\pm}_{\mathrm{train}}$ becomes biased towards the anomalous regions, reducing its ability to separate normal from anomalous samples (Qiu et al., 2022; Yoon et al., 2022). The existing literature examining the impact of contamination on unsupervised AD methods (Jiang et al., 2022; Qiu et al., 2022; Hien et al., 2024; Perini et al., 2023, 2022) typically considers contamination levels ranging from $0\%$ to $20\%$. Additionally, an analysis of 57 datasets spanning Natural Language Processing and Computer Vision in ADBench (Han et al., 2022) [Appendix B.2, Figure B1] revealed that nearly $70\%$ of the datasets exhibit anomaly ratios below 10%, with a median of $5\%$.
@@ 4 EPHAD: An evidence-based post-hoc adjustment framework
We consider the realistic scenario in which an AD model has already been trained on a possibly contaminated dataset $\mathcal{D}^{\pm}_{\mathrm{train}}$. Instead of retraining the model, our goal is to adapt its test-time predictions to mitigate the impact of contamination. To this end, we introduce a novel Evidence-based Post-Hoc Adjustment Framework for Anomaly Detection (EPHAD), that corrects model predictions using an evidence function at test-time. The evidence function $T(x):\mathcal{X}\to\mathbb{R}$ assigns higher values to samples deemed more likely to be normal and can incorporate domain-specific knowledge. Thus, EPHAD aligns with preparation-agnostic TTA methods (Xiao and Snoek, 2024).
@@
This objective balances two competing goals: aligning the adjusted PDF with the evidence (first term) and maintaining fidelity to the original PDF (second term). The temperature parameter $\beta$ controls this trade-off, recovering the evidence-driven solution as $\beta\to 0$ and reverting to the original model as $\beta\to\infty$. For the proof of (5), see Korbak et al. (2022). This interpretation also highlights a close connection to well-established TTA approaches used in generative models (Korbak et al., 2022; Mudgal et al., 2024; Li et al., 2024), where the model is viewed as an RL policy fine-tuned with a reward function encoding evidence or alignment criteria. In this view, EPHAD performs a KL-regularised shift of the contaminated density $f^{\pm}_{X}$ toward regions favored by the evidence function $T(x)$ while preserving consistency with $f^{\pm}_{X}$ through the KL term.